OpenStudy (anonymous):

A six-sided number cube with faces numbered 1,2,3,4,5 and 6 is rolled. What is the probability of rolling a 2 or a numbered greater than 5? A. 1/36 B. 1/3 C. 1/2 D.2/3

OpenStudy (anonymous):

They are asking what is the probability of rolling 2 or 6

OpenStudy (anonymous):

It is \[ \frac 1 6 + \frac 1 6 =\frac 1 3 \]

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

YW